# Retrieval-Oriented Knowledge for Click-Through Rate Prediction

### v1

@@ Abstract.
Click-through rate (CTR) prediction plays an important role in personalized recommendations. Recently, sample-level retrieval-based models (e.g., RIM) have achieved remarkable performance by retrieving and aggregating relevant samples. However, their inefficiency at the inference stage makes them impractical for industrial applications. To overcome this issue, this paper proposes a universal plug-and-play retrieval-oriented knowledge (ROK) framework. Specifically, a knowledge base, consisting of a retrieval-oriented embedding layer and a knowledge encoder, is designed to preserve and imitate the retrieved & aggregated representations in a decomposition-reconstruction paradigm. Knowledge distillation and contrastive learning methods are utilized to optimize the knowledge base, and the learned retrieval-enhanced representations can be integrated with arbitrary CTR models in both instance-wise and feature-wise manners. Extensive experiments on three large-scale datasets show that ROK achieves competitive performance with the retrieval-based CTR models while reserving superior inference efficiency and model compatibility.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: xx; Preprint; price: xisbn: xxx
@@ 1. Introduction
Click-through rate (CTR) prediction is a key component of many personalized online services, such as recommender systems (Xi et al., 2023) and web search (Lin et al., 2021; Fu et al., 2023; Dai et al., 2021). It seeks to estimate the probability of a user’s click given a particular context (Zhang et al., 2021). The CTR models could be mainly classified into two categories. The first category is the feature interaction-based methods. The core idea of these methods is to capture the high-order feature interactions across multiple fields with different operators (e.g., product (Qu et al., 2018; Wang et al., 2021; Guo et al., 2017; Chen et al., 2021b), convolution (Xin et al., 2019; Liu et al., 2019), and attention (Song et al., 2019; Xiao et al., 2017)). The second category is the user behavior modeling methods that leverage different architectures (e.g., RNN (Hidasi and Karatzoglou, 2018; Hidasi et al., 2015), CNN (Tang and Wang, 2018), attention (Zhou et al., 2019, 2018), memory bank (Pi et al., 2019; Ren et al., 2019)) to extract informative knowledge from user behavior sequences for final CTR prediction.
@@
To further enhance the performance, UBR4CTR (Qin et al., 2020) and SIM (Pi et al., 2020) retrieve useful behaviors from the user’s behavior history (i.e., clicked items), reducing the potential noise in user behavior sequences. Subsequent studies (Li et al., 2022b) have improved these methods’ efficiency using hashing functions (Chen et al., 2021a) and parallel retrieval execution (Cao et al., 2022) during inference. Recent studies (Qin et al., 2021; Du et al., 2022) further expand these retrieval-based methods from item-level retrieval to sample-level retrieval, applicable to general CTR prediction settings. Instead of retrieving similar items from the user history, RIM (Qin et al., 2021) adopts the idea of $k$ nearest neighbor ($k$NN) and designs a sample-centric retrieval method, which aggregates the relevant data samples retrieved from the search pool (e.g., the whole training dataset). PET (Du et al., 2022) constructs a hypergraph over the retrieved data instances and performs message propagation to get a better representation of the target data for final CTR prediction.
Although sample-level retrieval-based methods bring impressive performance enhancement, they have to perform instance-wise comparisons between the target data sample and each candidate sample in the search pool (usually million or even billion level). This leads to extreme inefficiency problems during inference, making it impractical for industrial applications, as illustrated in the left part of Figure 1. Recent work, such as DERT (Zheng et al., 2023), has improved the retrieval mechanisms by using vector retrieval techniques, enhancing efficiency. However, DERT’s dependency on the RIM Encoder and its $O(\log N)$ retrieval time complexity, along with the additional cost of encoding numerous samples in the retrieval pool, limits its scalability and effectiveness in large-scale applications.
@@
This work introduces the Retrieval-Oriented Knowledge (ROK) framework to tackle the inference inefficiency problem of sample-level retrieval-based methods, with two stages: Retrieval-Oriented Knowledge Construction and Knowledge Utilization. Initially, we pre-train a sample-level retrieval-based model, such as RIM(Qin et al., 2021). To resolve the inefficient inference issue, ROK tries to learn a neural network-based Knowledge Base to imitate the aggregated representations from the pre-trained retrieval-based methods via a Retrieval Imitation module, whose simple diagram is shown in Figure 1. The Knowledge Base, built on a novel decomposition-reconstruction paradigm, where a Retrieval-Oriented Embedding Layer captures the feature-wise embedding and a Knowledge Encoder reconstructs the instance-wise aggregated representations (i.e., retrieval-enhanced representation). In this way, instead of the time-consuming retrieval, efficient inference of neural networks could be adopted to get the approximated aggregated representations. Additionally, we introduce a Contrastive Regularization module to ensure learning stability and prevent model collapse. At the Knowledge Utilization stage, ROK designs two approaches (i.e., instance-wise and feature-wise) to integrate the retrieval-enhanced representations with various CTR models, thus providing high inference efficiency.
@@
  • In this work, we introduce ROK, a pioneering framework that transforms sample-level retrieval-based methods —previously deemed infeasible in the industry into a practical solution. Our approach innovatively leverages a neural network-based Knowledge Base to efficiently store the retrieval-enhance representations and hence obviates the need for time-intensive retrieval & aggregation during inference.
  • We use knowledge distillation and contrastive learning to optimizie the knowledge base, allowing the integration of retrieval-enhanced representations with various CTR models, both at the instance and feature levels.
  • Extensive experiments conducted across three large-scale datasets have shown that ROK not only achieves competitive performance when compared to existing retrieval-based CTR methods but also reserves superior inference efficiency. Moreover, ROK enhances the performance of various CTR methods due to its exceptional compatibility. This success underscores the effectiveness of the neural knowledge model as a compact surrogate for the retrieval pool.
@@
Figure 2. Overall framework of ROK. The process begins with the Retrieval-Oriented Knowledge Construction stage, where a sample-level retrieval-based model has already been pre-trained, and the retrieval and aggregation results are offline pre-stored. A tailored knowledge base is then designed to imitate the aggregation results, with Contrastive Regularization ensuring stabilization during the training phase. This design allows the time-consuming retrieval process to be substituted with the swift forward propagation of a neural network. In the subsequent Knowledge Utilization stage, the framework introduces the retrieval-enhanced representation into backbone CTR models in both an instance-wise and feature-wise approach, ensuring the maintenance of sample-level retrieval abilities while circumventing the complexities of retrieval and aggregation operations.
@@ 2.1. CTR Prediction
For the click-through rate (CTR) prediction, models can essentially be divided into two main categories: those focusing on feature interaction and those centering on user behavior modeling. The first category is the feature interaction-based methods, evolving from foundational works such as POLY2 (Chang et al., 2010) and Factorization Machines (FM) (Rendle, 2010). With the integration of Deep Neural Networks (DNNs), a variety of sophisticated deep feature interaction models have been proposed. These models aim to capture high-order feature interactions across different fields by employing various operations, such as the product operation(Qu et al., 2018; Wang et al., 2021; Guo et al., 2017; Chen et al., 2021b; Li et al., 2022a), convolution (Xin et al., 2019; Liu et al., 2019), and attention mechanisms (Song et al., 2019; Xiao et al., 2017). The key innovation of these models lies in their ability to identify and utilize complex interactions between a multitude of features to enhance the predictive performance of CTR models.
Besides, user behavior modeling is another core technique for CTR prediction that mines user preferences from historical interaction behaviors meticulously. To better extract informative knowledge from user’s behavior sequence, various network structures have been utilized, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks. GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both union-level and point-level. Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works have been proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). Among these, DIN and DIEN leverage the target-attention network to identify important historical items, while SASRec and BERT4Rec apply the self-attention network with Transformer architecture to excavate behavior dependencies. Besides, memory-based methods (Jiang et al., 2020) are also proposed to store user behavior representations in a read-write manner.
@@ 2.2. Retrieval-Augmented Recommendation
To further enhance the performance of CTR prediction, the retrieval-augmented recommendation is proposed, where the most relevant information is retrieved from the historical data. Specifically, UBR4CTR (Qin et al., 2020) and SIM (Ren et al., 2019) are designed to retrieve beneficial behaviors from the user’s historical extremely long behavior sequence, where UBR4CTR deploys the search engine method while SIM uses the hard search and soft search approaches. To make the search procedure end-to-end, ETA (Chen et al., 2022) is proposed by leveraging the SimHash algorithm to map the user behavior into a low-dimensional space, hence achieving learnable retrieval. Moreover, recent works further extend the retrieval-augmented recommendation from item-level retrieval to sample-level retrieval by retrieving informative samples. RIM (Qin et al., 2021) is the first to deploy this method that leverages the search engine to retrieve several relevant samples from the search pool and performs neighbor aggregation. PET (Du et al., 2022) and DERT (Zheng et al., 2023) have respectively made improvements in the interaction and retrieval mechanism of neighboring samples. PET constructs a hypergraph over the retrieved data samples and performs message propagation to improve the target data representations for final CTR prediction. DERT utilizes vector retrieval to speed up the retrieval process.
@@ 4.1. Overview of ROK
Despite the superior performance, sample-level retrieval-based methods suffer from the inference inefficiency problem due to the time-consuming online retrieval process. To this end, we propose a novel Retrieval-Oriented Knowledge (ROK) framework, where a knowledge base is built to imitate the aggregated retrieval knowledge $\mathbf{r}_{t}$. As shown in Figure 2, ROK consists of two stages: (1) Retrieval-Oriented Knowledge Construction, and (2) Knowledge Utilization.
In the first stage of ROK–retrieval-oriented knowledge construction–we pre-train a sample-level retrieval-based model [^2] In this work, we employed RIM(Qin et al., 2021) which is one of the foundational works in sample-level retrieval-based methods. The pre-training process exactly follows the original work. Importantly, our proposed method extends its applicability to encompass all sample-level retrieval-based methods., followed by designing a delicate knowledge base with a retrieval-oriented embedding layer and a knowledge encoder. This is designed to imitate the retrieval & aggregation results (i.e., $\mathbf{r}_{t}$) of the pre-trained model by directly generating the final aggregated representation $\mathbf{z}_{t}^{x}$ through the knowledge encoder. In this way, when online serving, the time-consuming retrieval process could be replaced by a simpler and faster forward propagation of a neural network (i.e., knowledge base). We adopt the mean square error loss for knowledge imitation and design a contrastive regularization loss to stabilize the learning process and prevent collapse (Jing et al., 2021; Shen et al., 2022).
@@ 4.2. Structure Design of Knowledge Base
For sample-level retrieval-based methods (Qin et al., 2021; Du et al., 2022), the overhead of inference inefficiency is mainly caused by the retrieval & aggregation operations, which are uncacheable and unavoidable during the online inference if a brand-new data sample comes. To this end, we propose a novel decomposition-reconstruction paradigm, where the aggregated representation $\mathbf{r}_{t}$ is first decomposed into feature-level embeddings, and then reconstructed via a neural encoder. As illustrated in  3, our proposed knowledge base comprises two key modules: (1) retrieval-oriented embedding layer $f$, and (2) knowledge encoder $g$.
Initially, we decompose the query $x_{t}$ into feature-wise embeddings via the retrieval-oriented embedding layer $f$. Then, the obtained feature-wise embeddings are fed into the knowledge encoder $g$ to reconstruct the instance-wise aggregated representation $\mathbf{z}_{t}^{x}=g(f(x_{t}))$. Specifically, a learnable embedding layer functions as the retrieval-oriented embedding layer. The architecture of knowledge encoder $g$ can be chosen arbitrarily (e.g., transformer (Vaswani et al., 2017)). For simplicity, we adopt multi-layer perceptron (MLP) in this paper. This approach reduces the time complexity of sample-level retrieval from $\mathcal{O}(N\log\left(N\right))$, where $N$ is the search pool size to $\mathcal{O}(1)$. Since we introduce the retrieval-oriented embedding layer $f$, we will cut down the embedding size of $f$ from $d$ to $d/2$ for fair comparison in space complexity. With the help of the learned knowledge base, online retrieval & aggregation operations can be avoided.

### v2

@@ Abstract.
Click-through rate (CTR) prediction is crucial for personalized online services. Sample-level retrieval-based models, such as RIM, have demonstrated remarkable performance. However, they face challenges including inference inefficiency and high resource consumption due to the retrieval process, which hinder their practical application in industrial settings. To address this, we propose a universal plug-and-play retrieval-oriented knowledge (ROK) framework that bypasses the real retrieval process. The framework features a knowledge base that preserves and imitates the retrieved & aggregated representations using a decomposition-reconstruction paradigm. Knowledge distillation and contrastive learning optimize the knowledge base, enabling the integration of retrieval-enhanced representations with various CTR models. Experiments on three large-scale datasets demonstrate ROK’s exceptional compatibility and performance, with the neural knowledge base serving as an effective surrogate for the retrieval pool. ROK surpasses the teacher model while maintaining superior inference efficiency and demonstrates the feasibility of distilling knowledge from non-parametric methods using a parametric approach. These results highlight ROK’s strong potential for real-world applications and its ability to transform retrieval-based methods into practical solutions. Our implementation code is available to support reproducibility [^1] https://github.com/HSLiu-Initial/ROK.git.
Information Retrieval; Recommender Systems; Knowledge Distillation; Contrastive Learning
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USA.booktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAisbn: 979-8-4007-0436-9/24/10doi: 10.1145/3627673.3679842ccs: Information systems Recommender systemsccs: Information systems Nearest-neighbor searchccs: Information systems Users and interactive retrieval
@@ 1. Introduction
Click-through rate (CTR) prediction is a key component of many personalized online services, such as recommender systems (Xi et al., 2023; Wang et al., 2023) and web search (Lin et al., 2021; Fu et al., 2023; Dai et al., 2021; Liu et al., 2024). It aims to estimate the probability of a user’s click given a particular context (Zhang et al., 2021). The CTR models can be mainly classified into two categories. The first category is the feature interaction-based methods. The core idea of these methods is to capture the high-order feature interactions across multiple fields with different operators (e.g., product (Qu et al., 2018; Wang et al., 2021; Guo et al., 2017; Chen et al., 2021b), convolution (Xin et al., 2019; Liu et al., 2019), and attention (Song et al., 2019; Xiao et al., 2017)). The second category is the user behavior modeling methods that leverage different architectures (e.g., RNN (Hidasi and Karatzoglou, 2018; Hidasi et al., 2015), CNN (Tang and Wang, 2018), attention (Zhou et al., 2019, 2018), memory bank (Pi et al., 2019; Ren et al., 2019)) to extract informative knowledge from user behavior sequences for final CTR prediction.
@@
Due to the problem of distribution shift and the daily generation of vast amounts of interaction data in industrial applications, using all the data for training becomes impractical. This results in a significant waste of historical data, as traditional methods cannot effectively utilize it. To address this issue, retrieval-based methods have been proposed. UBR4CTR (Qin et al., 2020) and SIM (Pi et al., 2020) retrieve useful behaviors from the user’s behavior history (i.e., clicked items), reducing the potential noise in user behavior sequences. Subsequent studies (Li et al., 2022b) improve these methods’ efficiency using hashing functions (Chen et al., 2021a) and parallel retrieval execution (Cao et al., 2022) during inference. Recent studies (Qin et al., 2021; Du et al., 2022) further expand these retrieval-based methods from item-level retrieval to sample-level retrieval, making them applicable to general CTR prediction settings. Instead of retrieving similar items from the user history, RIM (Qin et al., 2021) adopts the idea of $k$ nearest neighbor ($k$NN) and designs a sample-centric retrieval method, which aggregates the relevant data samples retrieved from the search pool (e.g., the whole training dataset).
Although sample-level retrieval-based methods brings impressive performance enhancement, they have to perform instance-wise comparisons between the target data sample and candidate samples in the search pool (usually at the million or even billion level), leading to extreme inefficiency problems during inference, as illustrated in the left part of Figure 1. While follow-up studies (Zheng et al., 2023; Li et al., 2024) have attempted to enhance retrieval mechanisms through vector retrieval techniques, they have instead created further challenges and failed to meet the necessary inference speed for deployment. Moreover, sample-level retrieval-based methods faces several other significant challenges. These include high resource consumption due to the linear increase in retrievers with request volume, as well as a divergent inference process that requires an inference-retrieval-inference cycle. This necessitates significant optimizations, which makes it challenging for industrial applications.
@@
This work introduces the Retrieval-Oriented Knowledge (ROK) framework to address the problems of sample-level retrieval-based methods by bypassing the real retrieval process. The framework operates in two stages: Retrieval-Oriented Knowledge Construction and Knowledge Utilization. In the first stage, we pre-train a sample-level retrieval-based method, such as RIM (Qin et al., 2021). To bypass the real retrieval process, ROK learns a neural network-based Knowledge Base to store the retrieval-oriented knowledge from this pre-trained method. Specifically, ROK imitates the aggregated representations from the pre-trained retrieval-based method via a Retrieval Imitation module, as illustrated in Figure 1. The Knowledge Base, built on a novel decomposition-reconstruction paradigm, where a Retrieval-Oriented Embedding Layer captures the feature-wise embedding and a Knowledge Encoder reconstructs the instance-wise aggregated representations (i.e., retrieval-enhanced representation). In this way, instead of the time-consuming retrieval, efficient inference of neural networks can be adopted to get the approximated aggregated representations. Additionally, we introduce a Contrastive Regularization module to ensure learning stability and prevent model collapse. At the Knowledge Utilization stage, ROK designs two approaches (i.e., instance-wise and feature-wise) to integrate the retrieval-enhanced representations with various CTR models, thus providing high inference efficiency.
@@
  • ROK is a novel framework that leverages knowledge distillation to efficiently utilize large amounts of historical data and retrieval-oriented knowledge for training and lightweight deployment. To the best of our knowledge, ROK is the first such framework. It transforms sample-level retrieval-based methods into a practical solution by using a neural network-based Knowledge Base to store retrieval-enhanced representations, eliminating the need for time-intensive retrieval during inference. This demonstrates the feasibility of distilling knowledge from non-parametric models and learning their inherent knowledge using a parametric approach.
  • ROK optimizes the Knowledge Base using a combination of knowledge distillation and contrastive learning. This approach allows for the seamless integration of retrieval-enhanced representations with various CTR models at both the instance and feature levels, enhancing their performance and compatibility.
  • Extensive experiments across three large-scale datasets demonstrate ROK’s exceptional compatibility and performance. The results show that ROK significantly enhances the performance of various CTR methods, indicating that the neural Knowledge Base serves as an effective and compact surrogate for the search pool. Although the knowledge distillation foundation initially suggests that ROK may only approach the teacher model’s performance, the introduction of contrastive regularization enables ROK to surpass the sample-level retrieval-based teacher model. This highlights ROK’s strong potential for real-world applications.
@@
Figure 2. Overall framework of ROK. The process begins with the Retrieval-Oriented Knowledge Construction stage, where a sample-level retrieval-based model has already been pre-trained, and the retrieval and aggregation results are offline pre-stored. A tailored knowledge base is then designed to imitate the aggregation results, with Contrastive Regularization ensuring stabilization during the training phase. This design allows the time-consuming retrieval & aggregration process to be substituted with the swift forward propagation of a neural network. In the subsequent Knowledge Utilization stage, the framework introduces the retrieval-enhanced representation into backbone CTR models in both an instance-wise and feature-wise approach, ensuring the maintenance of sample-level retrieval abilities while circumventing the complexities of retrieval & aggregation operations.
@@ 2.1. CTR Prediction
For click-through rate (CTR) prediction, models can essentially be divided into two main categories: those focusing on feature interaction and those centering on user behavior modeling. The first category is the feature interaction-based methods, evolving from foundational works such as POLY2 (Chang et al., 2010) and Factorization Machines (FM) (Rendle, 2010). With the integration of Deep Neural Networks (DNNs), a variety of sophisticated deep feature interaction models have been proposed. These models aim to capture high-order feature interactions across different fields by employing various operations, such as the product operation (Qu et al., 2018; Wang et al., 2021; Guo et al., 2017; Chen et al., 2021b; Li et al., 2022a), convolution (Xin et al., 2019; Liu et al., 2019), and attention mechanisms (Song et al., 2019; Xiao et al., 2017). The key innovation of these models lies in their ability to identify and utilize complex interactions between a multitude of features to enhance the predictive performance of CTR models.
Furthermore, user behavior modeling is another core technique for CTR prediction that mines user preferences from historical interaction behaviors meticulously. To better extract informative knowledge from a user’s behavior sequence, various network structures have been utilized, including Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), Attention Networks, and Memory Networks. GRU4Rec (Hidasi et al., 2015) designs Gated Recurrent Units (GRUs) to capture the preference-evolving relationship, while Caser (Tang and Wang, 2018) leverages the horizontal and vertical convolution to model skip behaviors at both the union-level and point-level. Moreover, the attention mechanism is the most popular method for modeling item dependencies, and several influential works have been proposed, including SASRec (Kang and McAuley, 2018), DIN (Zhou et al., 2018), DIEN (Zhou et al., 2019), and BERT4Rec (Sun et al., 2019). Among these, DIN and DIEN leverage the target-attention network to identify important historical items, while SASRec and BERT4Rec apply the self-attention network with Transformer architecture to excavate behavior dependencies. Additionally, memory-based methods (Jiang et al., 2020) are also proposed to store user behavior representations in a read-write manner.
@@ 2.2. Retrieval-Augmented Recommendation
To further enhance the performance of CTR prediction, retrieval-augmented recommendation is proposed, where the most relevant information is retrieved from historical data. Specifically, UBR4CTR (Qin et al., 2020) and SIM (Ren et al., 2019) are designed to retrieve beneficial behaviors from the user’s extremely long historical behavior sequence. UBR4CTR deploys the search engine method, while SIM uses hard search and soft search approaches. To make the search procedure end-to-end, ETA (Chen et al., 2022) is proposed by leveraging the SimHash algorithm to map user behavior into a low-dimensional space, hence achieving learnable retrieval. Moreover, recent works further extend retrieval-augmented recommendation from item-level retrieval to sample-level retrieval by retrieving informative samples. RIM (Qin et al., 2021) is the first to deploy this method, leveraging the search engine to retrieve several relevant samples from the search pool and perform neighbor aggregation. PET (Du et al., 2022) and DERT (Zheng et al., 2023) have respectively made improvements in the interaction and retrieval mechanisms of neighboring samples. PET constructs a hypergraph over the retrieved data samples and performs message propagation to improve the target data representations for final CTR prediction. DERT utilizes vector retrieval to speed up the retrieval process.
@@ 4.1. Overview of ROK
Despite their superior performance, sample-level retrieval-based methods suffer from challenges, including inference inefficiency and high resource consumption, due to the time-consuming online retrieval process. To this end, we propose a novel Retrieval-Oriented Knowledge (ROK) framework, where a knowledge base is built to imitate the aggregated retrieval knowledge $\mathbf{r}_{t}$. As shown in Figure 2, ROK consists of two stages: (1) Retrieval-Oriented Knowledge Construction, and (2) Knowledge Utilization.
In the first stage of ROK–retrieval-oriented knowledge construction–we pre-train a sample-level retrieval-based model [^3] In this work, as our research primarily focuses on inference efficiency, we choose a simple yet effective option that provides sufficiently good results: RIM (Qin et al., 2021), which is one of the foundational works in sample-level retrieval-based methods. The pre-training process strictly follows the original work (Qin et al., 2021). It is important to note that our proposed method can be generalized to all sample-level retrieval-based methods methods, extending its applicability and potential impact., followed by designing a delicate knowledge base with a retrieval-oriented embedding layer and a knowledge encoder. This is designed to imitate the retrieval & aggregation results (i.e., $\mathbf{r}_{t}$) of the pre-trained model by directly generating the final aggregated representation $\mathbf{z}_{t}^{x}$ through the knowledge encoder. In this way, when online serving, the time-consuming retrieval process could be replaced by a simpler and faster forward propagation of a neural network (i.e., knowledge base). We adopt the mean square error loss for knowledge imitation and design a contrastive regularization loss to stabilize the learning process and prevent collapse (Jing et al., 2021; Shen et al., 2022).
@@ 4.2. Structure Design of Knowledge Base
For sample-level retrieval-based methods (Qin et al., 2021; Du et al., 2022), the overhead of inference inefficiency and high resource consumption is mainly caused by the retrieval & aggregation operations, which are uncacheable and unavoidable during the online inference if a brand-new data sample comes. To this end, we manage to distill knowledge from non-parametric retrieval process using a parametric approach. To be specific, we propose a novel decomposition-reconstruction paradigm, where the aggregated representation $\mathbf{r}_{t}$ is first decomposed into feature-level embeddings, and then reconstructed via a neural encoder. As illustrated in  3, our proposed knowledge base comprises two key modules: (1) retrieval-oriented embedding layer $f$, and (2) knowledge encoder $g$.
Initially, we decompose the query $x_{t}$ into feature-wise embeddings via the retrieval-oriented embedding layer $f$. Then, the obtained feature-wise embeddings are fed into the knowledge encoder $g$ to reconstruct the instance-wise aggregated representation $\mathbf{z}_{t}^{x}=g(f(x_{t}))$. Specifically, a learnable embedding layer functions as the retrieval-oriented embedding layer. The architecture of knowledge encoder $g$ can be chosen arbitrarily (e.g., transformer (Vaswani et al., 2017)). For simplicity, we adopt multi-layer perceptron (MLP) in this paper. This approach reduces the time complexity of sample-level retrieval to $\mathcal{O}(1)$. Since we introduce the retrieval-oriented embedding layer $f$, we will cut down the embedding size of $f$ from $d$ to $d/2$ for fair comparison in space complexity. With the help of the learned knowledge base, online retrieval & aggregation operations can be avoided.